# SWAA: Sliding Window Attention Adaptation for Efficient Long-Context LLMs Without Pretraining

### v4

@@ Appendix A SWA’s Benefits and Each Method’s Drawbacks
SWA reduces the computational complexity to $O(N\cdot W)$, where $W$ is the window size. The benefits are threefold: (1) SWA reduces the computational load, (2) conserves GPU memory by limiting the required Key-Value (KV) cache, and (3) enhances KV cache reusability beyond traditional prefix caching, since a token’s state is independent of tokens outside its local window. However, there is no free lunch—each method of SWAA has some drawbacks, impairing the benefits brought by SWA to varying degrees.
FA Decode presents two primary drawbacks: (1) the benefits apply only to prefilling, while decoding speed is not accelerated as it utilizes full attention, and (2) the GPU memory required for the KV cache is not reduced, as the KV cache for the full context must be retained for decoding. In practice, however, many distributed LLM services have to recompute the KV cache of the entire chat history because storing and loading the KV cache complicates engineering systems, making prefilling occurs more frequently than expected, thereby amplifying the advantage of this method.
@@
CoT will greatly increase the generation length, especially for difficult tasks. So the decoding time will be much longer.
@@ Appendix B Some Problems of Other Long-context Benchmarks
    LongBench Bai et al. (2023) is classic and widely used, but its average context length (most are under 16k) is relatively short for modern models, i.e., it is already too easy. And its data source is too old, leading to a risk of test data leakage. So we choose not to use it.
@@
    RulerHsieh et al. (2024) has controllable context length, but its tasks are almost all synthetic and most of them are needle-retrieval tasks, thus failing to reflect the model’s overall long-context capability in real-world scenarios.
@@
    LongBench-V2 Bai et al. (2024b) is well-designed to necessitate deep understanding and reasoning over very long context. But it is too challenging for 4B-level models (e.g., Qwen3-4B-Thinking only gets 35% accuracy, which is too close to the random guessing baseline of 25%), making the improvement of different methods less distinguishable. Moreover, since it is in a multiple-choice question format, the results may not be sufficiently reliable because the model has a 25% chance of guessing the correct option.
    However, despite these drawbacks, they have been widely used in long-context model benchmarking. Thus we still elect to conduct our experiments on LongBench V2 Bai et al. (2024b) and Ruler Hsieh et al. (2024) to verify the generalizability of our conclusions, as shown in Appendix C.
Appendix C Results of Other Models
We show the results of Qwen3-30B-A3B in Table 4, and the results of Llama3.1-8B-Instruct in Table 5. Due to the time-intensive nature of training, we only test a selective set of configurations with fine-tuning, thus the results are fewer. For Llama3.1-8B-Instruct on Ruler, the context length is controlled to 32k due to its weaker long-context capability. From the results, we observe that for any benchmark, the trend of accuracy change is mostly consistent with our previous conclusions, demonstrating their generalizability.
@@ Appendix D Inference Efficiency
The TTFT, TPOT and total throughput when using vLLM are shown in Table 6. Since inference speed is highly dependent on hardware, implementation details, and workload characteristics, these numbers should be interpreted as reference values. However, from the results, we can still conclude that:
@@
In theory, FA Decode should yield a decoding speed identical to that of full attention. Yet, in this table, we observe acceleration on TPOT. This is because vLLM-v1 typically mixes different requests’ prefilling and decoding tokens in one sequence to improve GPU utilization. Thus, the speeds of prefilling and decoding may affect each other. If processing only a single request, the situation differs. For example, when the generation length is set to 2000, we find decoding takes over 95% of the total time, rendering the acceleration of the prefilling stage negligible—i.e., SWA with FA Decode is almost unable to improve efficiency in such cases.
@@ Appendix F Results of LightTransfer
LightTransfer Zhang et al. (2024) represents a promising attempt at SWA adaptation on full-attention models without pretraining. It proposes a layer selection method for SWA adaptation that calculates a "lazy ratio," represented by the ratio of attention from tokens at the end of the sequence (from a calibration dataset) to recent tokens versus global tokens. Layers with a higher "lazy ratio" are selected to apply SWA, while the rest retain full attention. This method is intuitive and theoretically sound, but our experiments reveal some negative results.
Since the complete code of LightTransfer is not open-source, we reproduce this method using LongAlign Bai et al. (2024a) as the calibration dataset for lazy layer detection, where the number of last tokens is set to 64, and the recent token window is set to 1024. From our experimental results shown in Table 8, we find that:
@@
Table 6: Efficiency metrics of different SWAA configurations on vLLM. "FA layers = 1/4" means one fourth of total layers use full attention while the others use SWA.
@@
Table 8: Results of LightTransfer on LongMemEval. "lazy" represents the half layers with higher lazy ratio, i.e. those which should apply SWA in theory. "non-lazy" represents the other part, i.e. those which should keep full attention.

### v5

@@ Appendix A SWA’s Benefits and Each Method’s Drawbacks
SWA reduces the computational complexity to $O(N\cdot W)$, where $W$ is the window size. The benefits are threefold: (1) SWA reduces the computational load, (2) conserves GPU memory by limiting the required Key-Value (KV) cache, and (3) enhances KV cache reusability beyond traditional prefix caching, since a token’s state is independent of tokens outside its local window. However, there is no free lunch: to preserve long-context capability, we have to use the methods of SWAA, each of which has drawbacks, impairing the aforementioned benefits to varying degrees.
FA Decode presents two primary drawbacks: (1) the benefits apply only to prefilling, while decoding speed is not accelerated as it utilizes full attention, and (2) the GPU memory required for the KV cache is not reduced, as the KV cache for the full context must be retained for decoding. Using CoT further increases the decoding time, reducing the speedup ratio. In practice, however, many distributed LLM services have to recompute the KV cache of the entire chat history because storing and loading the KV cache complicates engineering systems, making prefilling occurs more frequently than expected, thereby amplifying the advantage of this method.
@@ Appendix B Some Problems of Other Long-context Benchmarks
    LongBench (Bai et al., 2023) is classic and widely used, but its average context length (most are under 16k) is relatively short for modern models, i.e., it is already too easy. And its data source is too old, leading to a risk of test data leakage. So we choose not to use it.
@@
    Ruler (Hsieh et al., 2024) has controllable context length, but its tasks are almost all synthetic and most of them are needle-retrieval tasks, thus failing to reflect the model’s overall long-context capability in real-world scenarios.
@@
    LongBench-V2 (Bai et al., 2024b) is well-designed to necessitate deep understanding and reasoning over very long context. But it is too challenging for 4B-level models (e.g., Qwen3-4B-Thinking only gets 35% accuracy, which is too close to the random guessing baseline of 25%), making the improvement of different methods less distinguishable. Moreover, since it is in a multiple-choice question format, the results may not be sufficiently reliable because the model has a 25% chance of guessing the correct option.
    However, despite these drawbacks, they have been widely used in long-context model benchmarking. Thus we still elect to conduct our experiments on LongBench V2 (Bai et al., 2024b) and Ruler (Hsieh et al., 2024) to verify the generalizability of our conclusions, as shown in Appendix C.
Appendix C Experiment Results of Other Models
We show the accuracy results of Qwen3-30B-A3B in Table 3, and the results of Llama3.1-8B-Instruct in Table 4. Due to the time-intensive nature of training, we only test a selective set of configurations with fine-tuning. For Llama3.1-8B-Instruct on Ruler, the context length is controlled to 32k due to its weaker long-context capability. From the results, we observe that for any benchmark, the trend of accuracy change is mostly consistent with our previous conclusions, demonstrating their generalizability.
@@ Appendix D Inference Efficiency
The TTFT, TPOT and total throughput when using vLLM on a single H100 are shown in Table 5. Since inference speed is highly dependent on hardware, implementation details, and workload characteristics, these numbers should be interpreted as reference values. From the results, we can still conclude that:
@@
In theory, FA Decode should yield a decoding speed identical to that of full attention. Yet, in this table, we observe acceleration on TPOT. This is because vLLM-v1 typically mixes different requests’ prefilling and decoding tokens in one sequence to improve GPU utilization. Thus, the speeds of prefilling and decoding may affect each other. If processing only a single request, the situation differs. For example, when the generation length is set to 2000, we find decoding takes over 95% of the total time, rendering the acceleration of the prefilling stage negligible—i.e., SWA with FA Decode is almost unable to improve efficiency in such extreme cases.
@@ Appendix F Results of LightTransfer
LightTransfer (Zhang et al., 2024) represents a promising attempt at SWA adaptation on full-attention models without pretraining. It proposes a layer selection method for SWA adaptation that calculates a ”lazy ratio,” represented by the ratio of attention from tokens at the end of the sequence (from a calibration dataset) to recent tokens versus global tokens. Layers with a higher ”lazy ratio” are selected to apply SWA, while the rest retain full attention. This method is intuitive and theoretically sound, but our experiments reveal some negative results.
Since the complete code of LightTransfer is not open-source, we reproduce this method using LongAlign (Bai et al., 2024a) as the calibration dataset for lazy layer detection, where the number of last tokens is set to 64, and the recent token window is set to 1024. From our experimental results shown in Table 7, we find that:
@@
Table 5: Efficiency metrics of different SWAA configurations on vLLM. ”FA layers = 1/4” means one fourth of total layers use full attention while the others use SWA.
@@
Table 7: Results of LightTransfer on LongMemEval. ”lazy” represents the half layers with higher lazy ratio, i.e. those which should apply SWA in theory. ”non-lazy” represents the other part, i.e. those which should keep full attention.